OpenStudy (anonymous):

There are 20 students on the school's student council. A special homecoming dance committee is to be formed by randomly selecting 6 students from student council. How many possible committees can be formed? a) 720 b) 38,760 c) 27,907,200

OpenStudy (lgbasallote):

use this formula: \[\huge \frac{n!}{(n-r)!}\] \[\huge \implies \frac{20!}{(20-6)!}\] got it?

OpenStudy (anonymous):

I'm confused

OpenStudy (lgbasallote):

where?

OpenStudy (anonymous):

the whole thing.

OpenStudy (lgbasallote):

you have 20 students...that's why you have 20! 6 students are picked out of 20...that's why you get (20-6)!

OpenStudy (anonymous):

@lgbasallote, I believe this is a combination question.

OpenStudy (anonymous):

The order in which the students are chosen does not matter. So (a,b,c,d,e,f) is the same as (d,c,f,e,a,b).

OpenStudy (lgbasallote):

so it's \(\frac{n!}{r!(n-r)!}\)?

OpenStudy (anonymous):

@ashleys, you should use the formula \[C(n,m)=\frac{n!}{m!(n-m)!}.\]

OpenStudy (lgbasallote):

hmm okay..i get confused in these things...add a 6! there..the formula is \[\Huge \frac{n!}{r!(n-r)!}\] therefore you have \[\Huge \implies \frac{20!}{6!(20-6)!}\]

OpenStudy (anonymous):

I also get confused by these. I'm reading a post on them now.

OpenStudy (anonymous):

@ashleys, here is a really good post related to the last 6 or so questions you've posted: http://betterexplained.com/articles/easy-permutations-and-combinations/